Please help me figure this out.
Solve for x: 9^x+1 + 9^x+2 + 9^x+3 + 9^x+4 + 9^x+5 = 22,143
Answers:
a) -1/2
b) 1/2
c) -1/3
d) 1/3
e) 2/3

Answer by scott8148(6628)

factoring out 9^x+1 gives (9^x+1)(1+9+9^2+9^3+9^4)=22143

(9^x+1)(1+9+81+729+6561)=22143 __ (9^x+1)(7381)=22143 __ 9^x+1=3 9^x+1=9^(1/2)

x+1=1/2